# On growth of Sobolev norms for periodic non- linear Schrödinger and generalised Korteweg-de Vries equations under critical Gibbs dynamics

### v1

@@ Abstract.
We prove logarithmic growth bounds on Sobolev norms of the focusing mass-critical NLS and gKdV equations on the torus, which hold almost surely under the focusing Gibbs measure with optimal mass threshold constructed by Oh, Sosoe, and Tolomeo in [17]. More precisely, we will establish almost sure growth bounds for solutions $u(t)$ of the equations of the form
@@ 1. Introduction
(NLS) $\begin{cases}&i\partial_{t}u+\partial_{x}^{2}u+\absolutevalue{u}^{p-2}u=0\\ &u(0)=u_{0},\end{cases}$
@@
(1.2) $\mathrm{d}\rho_{p}(u)=\frac{1}{Z_{p}}e^{\frac{1}{p}\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{p}\,\mathrm{d}x}\,\mathrm{d}\mu_{0}(u).$
Here, $\mu_{0}$ is the Gaussian measure with inverse covariance $-\Delta$, formally given by
(1.3) $\mathrm{d}\mu_{0}(u)=\frac{1}{Z}e^{-\frac{1}{2}\int_{\operatorname{\mathbb{T}}}\absolutevalue{\partial_{x}u}^{2}\,\mathrm{d}x}\,\mathrm{d}u.$
@@
taking values in the space of periodic distributions $\mathcal{D}^{\prime}(\operatorname{\mathbb{T}})$, where $\{g_{n}\}_{n\in\operatorname{\mathbb{Z}}\setminus\{0\}}$ denotes a sequence of independent standard complex-valued Gaussian random variables. [^1] I.e. $\operatorname{Re}g_{n}$ and $\operatorname{Im}g_{n}$ are independent real-valued mean-zero Gaussian random variables with variance $\frac{1}{2}$. In the real-valued setting of (gKdV), we need to impose $g_{-n}=\overline{g_{n}}$ for all $n\in\operatorname{\mathbb{Z}}\setminus\{0\}$. An easy calculation shows that the measure $\mu_{0}$ is supported on $\bigcap_{s<\frac{1}{2}}H^{s}(\operatorname{\mathbb{T}})$. In the focusing case, $H$ is not bounded from below, and one can show that the density $e^{\frac{1}{p}\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{p}\,\mathrm{d}x}$ in (1.2) is not integrable with respect to $\mu_{0}$. For this reason, [14] considered the focusing Gibbs measure with a mass cutoff, [^2] In the defocusing case, $H$ is bounded from below and thus no mass cutoff is needed. i.e.
(1.4) $\mathrm{d}\rho_{p,K}(u)=\frac{1}{Z_{p,K}}e^{\frac{1}{p}\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{p}\,\mathrm{d}x}\mathbbm{1}_{\{M(u)\leq K\}}\,\mathrm{d}\mu_{0}(u),$
where $M(u)=\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{2}\,\mathrm{d}x$ denotes the mass and $K>0$. Recalling that the Hamiltonian $H$ and the mass $M$ are conserved under the NLS and gKdV dynamics, we expect (1.4) to be invariant measures whenever they exists.
We recall the normalisability of $\rho_{p,K}$; see [17] for details. Let $Q$ be the unique even and positive optimiser for the Gagliardo-Nirenberg-Sobolev inequality on $\operatorname{\mathbb{R}}$ (see Appendix B in [22] for details), namely
@@
at $p=6$, where $\norm{Q}_{L^{6}(\operatorname{\mathbb{R}})}^{6}=3\norm{\partial_{x}Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$. Then $\rho_{p,K}$ is normalisable (i.e. $Z_{p,K}<\infty$) exactly for $62<p<$ 6 and $K>0$, or $p=6$ and $K\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$. This was proven for the regimes $62<p<$ 6, $K>0$, and $p=6$, $K<\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$, as well as $p=6$, $K>\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$ in [14], although the proof contained a gap, and was corrected later in [4] for $p=6$ and $K>0$ sufficiently small. The proof of the normalisability and invariance at $p=6$ and $K<\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$ and even at the optimal threshold when $K=\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$ was completed in [17]. [^3] This result was somewhat unexpected due to the existence of minimal mass blow-up solutions $u_{*}$ to (NLS) with $\norm{u_{*}}_{L^{2}(\operatorname{\mathbb{T}})}=\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}$, constructed by Ogawa and Tsutsumi in [16]. There are also several works that study Gibbs measures for (NLS) in the infinite volume limit; see e.g. [18, 20, 23].
@@
(1.7) $\mathrm{d}\rho(u)=\frac{1}{Z_{K}}e^{\frac{1}{6}\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{6}\,\mathrm{d}x}\mathbbm{1}_{\{M(u)\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}\}}\,\mathrm{d}\mu(u).$
@@ 2. Bourgain’s strategy and a roadblock
Let $s>0$. Then (NLS) is locally well-posed in $H^{s}(\operatorname{\mathbb{T}})$. More precisely, there exists $\beta>0$ such that for all $M\geq 1$ and all $u_{0}\in H^{s}(\operatorname{\mathbb{T}})$ with $\norm{u_{0}}_{H^{s}(\operatorname{\mathbb{T}})}\leq M$ there exists a time $\tau=\tau(M)\gtrsim M^{-\beta}$ and a solution [^5] Uniqueness holds in a suitable $X^{s,b}$ space. $u\in C([-\tau,\tau];H^{s}(\operatorname{\mathbb{T}}))$ to (NLS) with $u|_{t=0}=u_{0}$, and $\sup_{t\in[-\tau,\tau]}\norm{u(t)}_{H^{s}(\operatorname{\mathbb{T}})}\leq 2M$. Moreover, the solution $u$ depends continuously on the initial data $u_{0}$.
Assume that we are given an invariant measure $\nu$ for (NLS) and assume that $\nu$ is absolutely continuous with respect to $\mu$ with Radon-Nikodym derivative $\frac{\mathrm{d}\nu}{\mathrm{d}\mu}$ in $L^{p}(\mu)$ for some $p>1$. We can then run Bourgain’s argument as follows: fixing $M\geq 1$ we have, by the local well-posedness Theorem 2, the invariance of $\nu$ under the (NLS) flow, and Hölder’s inequality, for any $s<\frac{1}{2}$
@@
$\leq\left(1+2\left\lfloor\frac{T}{\tau(M)}\right\rfloor\right)\norm{\frac{\mathrm{d}\nu}{\mathrm{d}\mu}}_{L^{p}(\mu)}\mu(\norm{u_{0}}_{H^{s}(\operatorname{\mathbb{T}})}\geq M)^{\frac{1}{p^{\prime}}}.$
Now, by a Gaussian tail estimate and again the local well-posedness Theorem 2, we obtain
$\nu\left(\sup_{t\in[-T,T]}\norm{u(t)}_{H^{s}(\operatorname{\mathbb{T}})}\geq 2M\right)\lesssim(1+2TM^{\beta})\norm{\frac{\mathrm{d}\rho}{\mathrm{d}\mu}}_{L^{p}(\mu)}e^{-\frac{M^{2}}{4p^{\prime}}}.$
@@
$\int_{\mathcal{D}^{\prime}(\operatorname{\mathbb{T}})}e^{\frac{p}{6}\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{6}\,\mathrm{d}x}\mathbbm{1}_{\{M(u)\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}\}}\,\mathrm{d}\mu(u)=\infty.$
@@
Identifying $\operatorname{\mathbb{T}}$ with $[-\frac{1}{2},\frac{1}{2})$, for some $\varepsilon>0$ and $\delta>0$ to be chosen later, take
@@
where $\chi\colon\operatorname{\mathbb{R}}\to[0,1]$ is a smooth cutoff function with support in $[-\frac{1}{4},\frac{1}{4}]$ and equal to $1$ on $[-\frac{1}{8},\frac{1}{8}]$. We have the following elementary estimates using smoothness and decay of $Q$.
@@
$\int_{\mathcal{D}^{\prime}(\operatorname{\mathbb{T}})}$ $e^{\frac{p}{6}\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{6}\,\mathrm{d}x}\mathbbm{1}_{\{M(u)\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}\}}\,\mathrm{d}\mu(u)$
$=\int_{\mathcal{D}^{\prime}(\operatorname{\mathbb{T}})}e^{\frac{p}{6}\lVert v-\widetilde{Q}_{\delta}^{\chi}\rVert_{L^{6}(\operatorname{\mathbb{T}})}^{6}}\mathbbm{1}_{\{M(v-\widetilde{Q}_{\delta}^{\chi})\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}\}}\,\mathrm{d}\mu\circ\tau_{\widetilde{Q}_{\delta}^{\chi}}(v)$
@@
$\int_{\mathcal{D}^{\prime}(\operatorname{\mathbb{T}})}$ $e^{\frac{p}{6}\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{6}\,\mathrm{d}x}\mathbbm{1}_{\{M(u)\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}\}}\,\mathrm{d}\mu(u)$
@@
$=(\widetilde{Q}_{\delta}^{\chi})^{6}+P(v,\widetilde{Q}_{\delta}^{\chi})-R(v,\widetilde{Q}_{\delta}^{\chi}),$
@@
$\int_{\mathcal{D}^{\prime}(\operatorname{\mathbb{T}})}$ $e^{\frac{p}{6}\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{6}\,\mathrm{d}x}\mathbbm{1}_{\{M(u)\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}\}}\,\mathrm{d}\mu(u)$
@@
where in the last line, we used the fact that $v\mapsto R(v,\widetilde{Q}_{\delta}^{\chi})$ is odd and $\mu$ is symmetric about 00, and that the event $\{\norm{v}_{L^{2}(\operatorname{\mathbb{T}})}\leq\frac{\varepsilon}{2}\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}\}$ has some positive probability $\varepsilon_{1}$. Appealing to Lemma 4, we finally obtain
$\int_{\mathcal{D}^{\prime}(\operatorname{\mathbb{T}})}$ $e^{\frac{p}{6}\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{6}\,\mathrm{d}x}\mathbbm{1}_{\{M(u)\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}\}}\,\mathrm{d}\mu(u)$
@@ 3. Orlicz spaces and the construction of a Young function
We call a function $\Psi\colon\operatorname{\mathbb{R}}\to[0,\infty)$ a Young function if it is an even, continuous, increasing, convex function with $\Psi(0)=0$ and $\lim_{x\to\infty}\frac{\Psi(x)}{x}=\infty$.
@@
One can check that after identifying functions that are equal $\mu$-almost everywhere, $\norm{\cdot}_{F}$ defines a norm. Note that $\Psi\colon x\mapsto|x|^{p}$ returns the spaces $\Psi(L)=L^{p}$. Importantly, Orlicz spaces enjoy a Hölder-type inequality: for all $f\in\Psi(L)$ and all $g\in\Psi^{*}(L)$ we have
(3.1) $\int_{\Omega}\absolutevalue{fg}\,\mathrm{d}\nu\leq 2\norm{f}_{\Psi(L)}\norm{g}_{\Psi^{*}(L)}.$
@@
for $x>0$, where $q$ and $10<\alpha<$ 1 are chosen later, and $F(0)\coloneqq 0$.
@@
First, $F(0)=0$, $F$ is even, continuous, increasing and $\lim_{x\to\infty}\frac{F(x)}{x}=\infty$. To check that $F$ is convex we compute for $x>1$
@@
This shows that $G$ is convex on $(e^{1-\alpha},\infty)$ and hence that $F$ is convex on $\operatorname{\mathbb{R}}$. ∎
@@ 4. Our strategy
Choosing $C$ such that $-cC<-2$, we have
@@
and where we used the fact that $F^{*}(0)=0$, $F^{*}$ is increasing, the tail estimate (from which we have the constant $c_{s}$), and Lemma 8. Here $M_{0}$ is such that
@@
Since $10<\alpha<$ 1 there exists $c_{1}=c_{1}(\alpha,q,s)$ such that
@@ 5.2. Stability of GNS optimisers
$\begin{split}S_{\gamma}=\big{\{}u\in L^{2}(\operatorname{\mathbb{T}})&:M(u)\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2},\\ &\norm{\mathbf{P}_{\leq k}u_{\neq 0}}_{L^{6}(\operatorname{\mathbb{T}})}^{6}\leq(C_{\rm GNS}-\gamma)\norm{\partial_{x}\mathbf{P}_{\leq k}u}_{L^{2}(\operatorname{\mathbb{T}})}^{2}\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{4}\text{ for }k\geq 1\big{\}}.\end{split}$
@@
(5.4) $\int_{S_{\gamma}}e^{\frac{r}{6}\int_{\operatorname{\mathbb{T}}}\absolutevalue{u}^{6}\,\mathrm{d}x}\,\mathrm{d}\mu(u)<\infty.$
@@ 5.3. A change-of-variable formula
The following proposition shows that a small neighbourhood of $\mathcal{M}_{\delta^{\ast}}$ can be endowed with an orthogonal coordinate system in terms of $e^{i\theta}\partial_{\delta}Q_{\delta,x_{0}}^{\chi}$, $e^{i\theta}\partial_{x_{0}}Q_{\delta,x_{0}}^{\chi}$, $ie^{i\theta}Q_{\delta,x_{0}}^{\chi}$ and $V_{\delta,x_{0},\theta}$. Recalling that $\mu$ is supported on functions $u\in H^{s}(\operatorname{\mathbb{T}})$ for $s<\frac{1}{2}$, we redefine $U_{\varepsilon,\delta^{*}}$ to mean $U_{\varepsilon,\delta^{*}}\cap(H^{\frac{1}{2}})^{c}$. Next, given a small $\delta>0$ and $x_{0}\in\operatorname{\mathbb{T}}$ we let $V_{\delta,x_{0},0}=V_{\delta,x_{0},0}\cap H^{1}(\operatorname{\mathbb{T}})$. Let $\mu_{\delta,x_{0}}^{\perp}$ denote the Gaussian measure with $V_{\delta,x_{0},0}\subset H^{1}(\operatorname{\mathbb{T}})$ as its Cameron-Martin space. Then we have the following change of variables for the $\mu$-integration over $U_{\varepsilon,\delta^{*}}$.
@@
[…]
@@
$J(v)=(2\delta^{-2}-1)\langle Q_{\delta}^{\chi},\operatorname{Re}v\rangle_{L^{2}(\operatorname{\mathbb{T}})}+(1+\eta)\langle(Q_{\delta}^{\chi})^{4},2(\operatorname{Re}v)^{2}+\frac{1}{2}\absolutevalue{v}^{2}\rangle_{L^{2}(\operatorname{\mathbb{T}})}.$
@@
$\int_{\{\norm{v}_{L^{2}(\operatorname{\mathbb{T}})}\leq\varepsilon_{1}\}}e^{C^{\prime}_{\eta}\int_{\operatorname{\mathbb{T}}}\absolutevalue{v}^{6}\,\mathrm{d}x}\,\mathrm{d}\mu^{\perp}_{\delta}(v)<\infty$

### v2

@@ Abstract.
We prove logarithmic growth bounds on Sobolev norms of the focusing mass-critical NLS and gKdV equations on the torus, which hold almost surely under the focusing Gibbs measure with optimal mass threshold constructed by Oh, Sosoe, and Tolomeo in [Invent. Math. 227 (2022), no. 3, 1323–1429]. More precisely, we will establish almost sure growth bounds for solutions $u(t)$ of the equations of the form
@@ 1. Introduction
[…]
@@
[…]
Here, $\mu_{0}$ is the Gaussian measure with inverse covariance $\Delta$, formally given by
[…]
@@
[…]
@@
at $p=6$, where $\norm{Q}_{L^{6}(\operatorname{\mathbb{R}})}^{6}=3\norm{\partial_{x}Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$. Then $\rho_{p,K}$ is normalisable (i.e. $Z_{p,K}<\infty$) exactly for $2<p<6$ and $K>0$, or $p=6$ and $K\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$. This was proven for the regimes $2<p<6$, $K>0$, and $p=6$, $K<\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$, as well as $p=6$, $K>\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$ in [14], although the proof contained a gap, and was corrected later in [4] for $p=6$ and $K>0$ sufficiently small. The proof of the normalisability and invariance at $p=6$ and $K<\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$ and even at the optimal threshold when $K=\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}$ was completed in [17]. [^3] This result was somewhat unexpected due to the existence of minimal mass blow-up solutions $u_{*}$ to (NLS) with $\norm{u_{*}}_{L^{2}(\operatorname{\mathbb{T}})}=\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}$, constructed by Ogawa and Tsutsumi in [16]. There are also several works that study Gibbs measures for (NLS) in the infinite volume limit; see e.g. [18, 20, 23].
@@
[…]
@@ 2. Bourgain’s strategy and a roadblock
Let $s>0$. Then (NLS) is locally well-posed in $H^{s}(\operatorname{\mathbb{T}})$. More precisely, there exists $\beta>0$ such that for all $M\geq 1$ and all $u_{0}\in H^{s}(\operatorname{\mathbb{T}})$ with $\norm{u_{0}}_{H^{s}(\operatorname{\mathbb{T}})}\leq M$ there exists a time $\tau=\tau(M)\gtrsim M^{-\beta}$ and a solution [^5] Uniqueness holds in a suitable $X^{s,b}$ space. $u\in C([-\tau,\tau];H^{s}(\operatorname{\mathbb{T}}))$ to (NLS) with $u(0)=u_{0}$, and $\sup_{t\in[-\tau,\tau]}\norm{u(t)}_{H^{s}(\operatorname{\mathbb{T}})}\leq 2M$. Moreover, the solution $u$ depends continuously on the initial data $u_{0}$.
Assume that we are given an invariant measure $\nu$ for (NLS) and assume that $\nu$ is absolutely continuous with respect to $\mu$ with Radon-Nikodym derivative $\frac{\mathrm{d}\nu}{\mathrm{d}\mu}$ in $L^{p}(\mu)$ for some $p>1$. We can then run Bourgain’s argument as follows: fixing $M\geq 1$ we have, by the local well-posedness Theorem 2, the invariance of $\nu$ under the (NLS) flow, and Hölder’s inequality,
@@
$\leq\left(1+2\left\lfloor\frac{T}{\tau(M)}\right\rfloor\right)\norm{\frac{\mathrm{d}\nu}{\mathrm{d}\mu}}_{L^{p}(\mu)}\mu(\norm{u_{0}}_{H^{s}(\operatorname{\mathbb{T}})}\geq M)^{\frac{1}{p^{\prime}}}$
for any $s<\frac{1}{2}$. Now, by a Gaussian tail estimate and again the local well-posedness Theorem 2, we obtain
$\nu\left(\sup_{t\in[-T,T]}\norm{u(t)}_{H^{s}(\operatorname{\mathbb{T}})}\geq 2M\right)\lesssim(1+2TM^{\beta})\norm{\frac{\mathrm{d}\nu}{\mathrm{d}\mu}}_{L^{p}(\mu)}e^{-\frac{M^{2}}{4p^{\prime}}}.$
@@
[…]
@@
Identifying $\operatorname{\mathbb{T}}$ with $\frac{1}{2},\frac{1}{2})$, for some small $\varepsilon>0$ and $\delta>0$ to be chosen later, take
@@
where $\chi\colon\operatorname{\mathbb{R}}\to[0,1]$ is a radially-decreasing smooth cutoff function with support in $\frac{1}{4},\frac{1}{4}]$ and equal to $1$ on $\frac{1}{8},\frac{1}{8}]$. We have the following elementary estimates using smoothness and exponential decay of $Q$ and $\partial_{x}Q$ (see Proposition B.7 in [22]).
@@
[…]
$=\int_{\mathcal{D}^{\prime}(\operatorname{\mathbb{T}})}e^{\frac{p}{6}\lVert v-\widetilde{Q}_{\delta}^{\chi}\rVert_{L^{6}(\operatorname{\mathbb{T}})}^{6}}\mathbbm{1}_{\{M(v-\widetilde{Q}_{\delta}^{\chi})\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2}\}}\,\mathrm{d}\mu\circ\tau_{-\widetilde{Q}_{\delta}^{\chi}}(v)$
@@
[…]
@@
$=\big\lVert\widetilde{Q}_{\delta}^{\chi}\big\rVert_{L^{6}(\operatorname{\mathbb{T}})}^{6}+P(v,\widetilde{Q}_{\delta}^{\chi})-R(v,\widetilde{Q}_{\delta}^{\chi}),$
@@
[…]
@@
where in the last line, we used the fact that $v\mapsto R(v,\widetilde{Q}_{\delta}^{\chi})$ is odd and $\mu$ is symmetric about 0, and that the event $\{\norm{v}_{L^{2}(\operatorname{\mathbb{T}})}\leq\frac{\varepsilon}{2}\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}\}$ has some positive probability $\varepsilon_{1}$. Appealing to Lemma 4, we finally obtain
[…]
@@ 3. Orlicz spaces and the construction of a Young function
We call a function $\Psi\colon\operatorname{\mathbb{R}}\to[0,\infty)$ a Young function if it is an even, continuous, increasing on $0,\infty)$, convex function with $\Psi(0)=0$ and $\lim_{x\to\infty}\frac{\Psi(x)}{x}=\infty$.
@@
One can check that after identifying functions that are equal $\mu$-almost everywhere, $\norm{\cdot}_{\Psi(L)}$ defines a norm. Note that $\Psi\colon x\mapsto|x|^{p}$ returns the spaces $\Psi(L)=L^{p}$. Importantly, Orlicz spaces enjoy a Hölder-type inequality: for all $f\in\Psi(L)$ and all $g\in\Psi^{*}(L)$ we have
[…]
@@
for $x>0$, where $q>0$ and $0<\alpha<1$ are chosen later, and $F(0)\coloneqq 0$.
@@
It is easy to see that $F(0)=0$, $F$ is even, continuous, increasing on $0,\infty)$ and $\lim_{x\to\infty}\frac{F(x)}{x}=\infty$. To check that $F$ is convex we compute for $x>1$
@@
[…]
@@ 4. Our strategy
Choosing $C$ such that cC<2-cC<-2, we have
@@
and where we used the fact that $^{*}(0)=0, FF^{*}$ is increasing, the tail estimate (from which we have the constant $c_{s}$), and Lemma 8. Here $M_{0}$ is such that
@@
Since $0<\alpha<1$ there exists $c_{1}=c_{1}(\alpha,q,s)$ such that
@@ 5.2. Stability of GNS optimisers
$\begin{split}S_{\gamma}=\big\{u\in L^{2}&(\operatorname{\mathbb{T}}):M(u)\leq\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{2},\\ &\norm{\mathbf{P}_{\leq k}u_{\neq 0}}_{L^{6}(\operatorname{\mathbb{T}})}^{6}\leq(C_{\rm GNS}-\gamma)\norm{\partial_{x}\mathbf{P}_{\leq k}u}_{L^{2}(\operatorname{\mathbb{T}})}^{2}\norm{Q}_{L^{2}(\operatorname{\mathbb{R}})}^{4}\text{ for all }k\geq 1\big\}.\end{split}$
@@
[…]
@@ 5.3. A change-of-variable formula
The following proposition shows that a small neighbourhood of $\mathcal{M}_{\delta^{\ast}}$ can be endowed with an orthogonal coordinate system in terms of $e^{i\theta}\partial_{\delta}Q_{\delta,x_{0}}^{\chi}$, $e^{i\theta}\partial_{x_{0}}Q_{\delta,x_{0}}^{\chi}$, $ie^{i\theta}Q_{\delta,x_{0}}^{\chi}$ and $V_{\delta,x_{0},\theta}$. Recalling that $\mu$ is supported on functions $u\in H^{s}(\operatorname{\mathbb{T}})$ for $s<\frac{1}{2}$ and that $\mu(H^{\frac{1}{2}})=0$, we redefine $U_{\varepsilon,\delta^{*}}$ to mean $U_{\varepsilon,\delta^{*}}\cap(H^{\frac{1}{2}})^{c}$. Next, given a small $\delta>0$ and $x_{0}\in\operatorname{\mathbb{T}}$ we let $V_{\delta,x_{0},0}=V_{\delta,x_{0},0}\cap H^{1}(\operatorname{\mathbb{T}})$. Let $\mu_{\delta,x_{0}}^{\perp}$ denote the Gaussian measure with $V_{\delta,x_{0},0}\subset H^{1}(\operatorname{\mathbb{T}})$ as its Cameron-Martin space. Then we have the following change of variables for the $\mu$-integration over $U_{\varepsilon,\delta^{*}}$.
@@
Fix $K>0$ and sufficiently small $\delta^{\ast}>0$. [^8] The choice of $\delta^{\ast}$ comes from Proposition 6.4 in [17], which introduces an orthogonal coordinate system for $U_{\varepsilon,\delta^{\ast}}$. Let $F(u)\geq 0$ be a functional of $u$ on $\operatorname{\mathbb{T}}$ which is continuous in some topology $H^{s}(\operatorname{\mathbb{T}})$, $s<\frac{1}{2}$, such that $F\leq C$ for some $C$ and $F(u)=0$ if $M(u)>K$. Then there is a locally finite measure $\sigma$ on $(0,\delta^{\ast})$ such that
@@
[…]
@@
[…]